OpenStudy (anonymous):

PLEASE give me guidance? What is the value of y in the solution to the following system of equations?5x - 3y = -32x - 6y = -6 Answers: -1 0 1 2

OpenStudy (anonymous):

can u draw the graphs? thats very conceptual

OpenStudy (anonymous):

5x - 3y = -32 ----- A x - 6y = -6 ------- B from equation B x - 6y = -6 => x= 6y-6 put this in eaquation A and solve for y

OpenStudy (anonymous):

Okay, thank you, I'm trying

OpenStudy (anonymous):

u r welcome

OpenStudy (anonymous):

Question!

OpenStudy (anonymous):

I see that you got x - 6y = -6 => x= 6y-6, but how did you get from "x - 6y, to x= 6y?" Because we're solving for "x"?

OpenStudy (anonymous):

Nevermind, I found my answer to that, but I still don't know how to go about solving for "X"